OpenStudy (anonymous):

solve and graph -3/4 > and equal to w + 1

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathmale

hero (hero):

Hint: Subtract 1 from both sides to isolate w

OpenStudy (mathmale):

Panda: I'm delighted to work with you. But there's a "however:" I wish you would try to apply some of what you learned in the previous problem to solving the current problem BEFORE you ask all these people for help.

OpenStudy (anonymous):

I SUBTRACTED and that gives me 1 3/4>w

OpenStudy (mathmale):

\[-3/4 \ge w + 1\] can be written in slightly clearer form as

OpenStudy (mathmale):

Please, try to develop some faith in your own ability to solve problems. At least give each problem a try before you ask for help. This might actually be faster in the long run.

OpenStudy (anonymous):

Is the answer 7/4 > w

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

i got the answer by subtracting 1 from both

OpenStudy (anonymous):

sides

OpenStudy (mathmale):

\[-3/4 \ge w + 1\] .... if you now subtract 1 from both sides, you get\[-(3/4) - 1 \ge w + 1-1, ~or ~-(3/4) - 1 \ge w \] Please simplify this.

OpenStudy (mathmale):

Combine:\[\frac{ -3 }{ 4 }-1\]

OpenStudy (anonymous):

3/4

OpenStudy (mathmale):

Sorry, but no. Try this: 3 - --- 4 + 1 ---------

OpenStudy (mathmale):

Or use your calculator: -0.75 -1.00 -------- ?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

but both quantities have the same sign, so combining them must produce a result with the same sign. That sign is (-), not (+).

OpenStudy (mathmale):

Please do this : combine 3/4 and 1. What do you get?

OpenStudy (anonymous):

1/4

OpenStudy (mathmale):

You are subtracting, Panda. I've asked you to ADD. Try again. Have you a calculator? Do this both by hand and by calculator.

OpenStudy (mathmale):

If I type out: (3/4) + 1, that automatically asks you to "ADD."

OpenStudy (mathmale):

If, on the other hand, I type out -(3/4) - 1, THAT AUtomatically asks you to ADD again, BUT the result will have a (-) sign.

OpenStudy (mathmale):

So, y ou have: -(3/4) - 1 EQUAL TO OR GREATER THAN w. Combine the -(3/4) and the -1 now, please.

OpenStudy (mathmale):

1 + (3/4) = 1 3/4, or 1.75, or 7/4. -1 - (3/4) = -1.75, or -7/4 So your inequality becomes \[\frac{ -7 }{ 4 }\ge w\]

OpenStudy (mathmale):

or \[w \le -7/4\]

OpenStudy (mathmale):

Haven't heard from you for a while. OpenStudy says you are "just looking around." The practice y ou're getting here is really important. Next time, would you please stay with the problem until completion? or excuse yourself if you're tired and want a break? Thanks.